OpenStudy (anonymous):

The perimeter of a rectangle with base 10 in. and height 13 in. is _____?

Directrix (directrix):

2(10) + 2 ( 13) = 20 + 26 = 46 inches

Directrix (directrix):

Opposite sides of a rectangle are congruent.